OpenStudy (anonymous):

bayside insurance offers two health plans. Under plan A, Giselle would have to pay the first $80 of her medical bills, plus 40% of the rest. Under plan B, Giselle would pay the first $100, but only 30% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $100 in bills. Giselle would save with plan B if she had more than $--- in bills. round to the nearest whole number.

OpenStudy (amistre64):

a(x) = 80 + .4x b(x) = 100 +.3x

OpenStudy (amistre64):

when they equal we have an equilibrium and anything past it flips the money

OpenStudy (amistre64):

100 +.3x = 80 + .4x ; -.3x and -80 20 = .1x x = 20/.1 = 200/1 = 200

OpenStudy (amistre64):

plan b saves money for anything over 200 dollars

OpenStudy (anonymous):

Hi amistre64 i posted the answer and 200 is not correct

OpenStudy (amistre64):

then i probably left out something ....

OpenStudy (anonymous):

would you double check for me please if its not asking too much

OpenStudy (amistre64):

i will, but make sure you are apllying the right format to your answer in the program your using too

OpenStudy (anonymous):

This is the question at the end that needs the answer Giselle would save with plan B if she had more than $--- in bills. round to the nearest whole number.

OpenStudy (amistre64):

well; when the bill costs 50 lets say; then she pays out of pocket 50 dollars..... and if we plug in a 50 here it doesnt say 50...so i prolly missed something

OpenStudy (anonymous):

so 50 is the answer

OpenStudy (amistre64):

50 was just a test

OpenStudy (amistre64):

the bill comes in; and its say 500 500 - 80 = whats left .4 (500-80) = what she owes for it; .4(530) = 212; so she pays a total of .4(B-80) + 80

OpenStudy (amistre64):

lets test that with 400 dollar bill .4(400-80) + 80 = Payment of 208 out of pocket looks good

OpenStudy (amistre64):

if the bill = 80 then; we get .4(0) + 80 = 80 out of pocket

OpenStudy (amistre64):

.4B -32+80 = .4x +48 for plan a then right?

OpenStudy (amistre64):

.3(x-100)+100 .3x -30+100 .3x +70 for plan b

OpenStudy (amistre64):

now we can equate them .4x +48 = .3x +70 .1x = 22 x = 220 .... maybe?

OpenStudy (anonymous):

hot dog amistre64 you're awesome that is correct thank you you are a life saver

OpenStudy (amistre64):

:) youre welcome